Myrick Woodcraft Company (MWC) manufactures “antique” wooden cabinets to house modern televisions. MWC began operations in January of last year. Sidney Myrick, the owner, asks for your assistance. He believes that he needs to better understand the cost of the cabinets for pricing purposes. You have collected the following data concerning actual production over the past year:

Month Number of Cabinets Produced Total Cost
January 800 $ 21,000
February 3,600 32,500
March 1,960 29,500
April 600 18,600
May 1,600 29,000
June 1,300 27,000
July 1,100 25,600
August 1,800 31,000
September 2,280 32,000
October 2,940 31,500
November 3,280 32,000
December 400 16,500


Required

  1. To understand the department’s cost behavior, you decide to plot the points on graph paper and sketch a total cost line.

  1. (1) Enter the number of units and their costs in increasing order.

  1. Using the high-low method, compute the total cost equation for the preceding data.

  1. (1) Compute the variable cost per unit.

  2. (2) Compute total fixed costs.

  3. (4) Calculate the total cost assuming 2,000 cabinets are made.

Myrick Woodcraft Company (MWC) manufactures “antique” wooden cabinets to house modern televisions. MWC began operations in January of last year. Sidney Myrick, the owner, asks for your assistance. He believes that he needs to better understand the cost of the cabinets for pricing purposes. You have collected the following data concerning actual production over the past year:

Month Number of Cabinets Produced Total Cost
January 800 $ 21,000
February 3,600 32,500
March 1,960 29,500
April 600 18,600
May 1,600 29,000
June 1,300 27,000
July 1,100 25,600
August 1,800 31,000
September 2,280 32,000
October 2,940 31,500
November 3,280 32,000
December 400 16,500


Required

  1. To understand the department’s cost behavior, you decide to plot the points on graph paper and sketch a total cost line.

  1. (1) Enter the number of units and their costs in increasing order.

  1. Using the high-low method, compute the total cost equation for the preceding data.

  1. (1) Compute the variable cost per unit.

  2. (2) Compute total fixed costs.

  3. (4) Calculate the total cost assuming 2,000 cabinets are made.
Answer #1

Solution :

1. Statement showing number of units and their costs in an increasing order

Number of Cabinets Produced Total Cost ( $ )
400 16500
600 18600
800 21000
1100 25600
1300 27000
1600 29000
1800 31000
1960 29500
2280 32000
2940 31500
3280 32000
3600 32500

1.

Computation of the total cost equation applying High-Low Method, (Refer Working Note 1 & 2)

Total Cost Equation = Variable Cost + Fixed Cost

= Number of units produced * 5 + 14500

Working Note 1:

Variable cost per unit = (Highest Activity Cost - Lowest Activity Cost) / (Highest Activity Units - Lowest Activity Units)

= (32500 - 16500) / (3600 - 400)

= 16000 / 3200

= $5 per unit

Working Note 2:

Fixed Cost = Highest Activity Cost - (Variable Cost per unit * Highest Activity Units)

= 32500 - (5*3600)

= $14,500

4) Calculation of the total cost assuming 2,000 cabinets are made

Total Cost (2,000 cabinets) = Total Variable cost + Total Fixed Cost)

= 2000*5 + 14500

= 10000+14500

= $24,500